OpenStudy (anonymous):

what is the common difference of a 57- arithmetic sequence where the 1st term is -25 and he sum is 17,727 ?

OpenStudy (anonymous):

\[\LARGE S_{57}=\frac {57}{2}[2\cdot a_1+(57-1)d]\] \[\LARGE 17727=\frac{57}{2}[2\cdot (-25)+56d]\] find d (difference.)

OpenStudy (anonymous):

d= 14535 ?

OpenStudy (anonymous):

it can't be lol ... then a_1=-25 a_2=14515 a_3 ... you have a_57=17727 that's impossible... try again :)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

where are you confused...can you show how did you come up with d=14535 ? :)

OpenStudy (anonymous):

im confused on how to get d

OpenStudy (anonymous):

ok ... but do you mind showing what you've worked so far, so I can see where you got it wrong. ?

OpenStudy (anonymous):

57/2 ( 2. a1 (56) d this is where i get confused

OpenStudy (anonymous):

you have a1=-25 \[\LARGE 17727=\frac{57}{2}[-50+56d]\] now multiply both sides by 2 and you'll get: \[\LARGE 2\cdot 17727=57[-50+56d]\] now divide both sides by 57 and you'll get: \[\LARGE \frac{2\cdot 17727}{57}=-50+56d\] now grab a calculator and calculate the left side and tell me what you get... so we can go further . ;)

OpenStudy (anonymous):

622

OpenStudy (anonymous):

d is 12

OpenStudy (anonymous):

ok/// we have: \[\LARGE 622=-50+56d\] now add 50 to both sides 622+50=-50+50+56d 672=56d now divide both sides by 56 and tell me what you get

OpenStudy (anonymous):

yes d=12 well done...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

My pleasure :)